OpenStudy (sugarapples):

√ 24 can I get help please

jimthompson5910 (jim_thompson5910):

can you think of a factor of 24 that is a perfect square?

OpenStudy (sugarapples):

6 and 4

jimthompson5910 (jim_thompson5910):

6 isn't a perfect square, but 4 is

jimthompson5910 (jim_thompson5910):

you're on the right track though 24 = 4*6 so \[\Large \sqrt{24} = \sqrt{4*6}\] \[\Large \sqrt{24} = \sqrt{4}*\sqrt{6}\] \[\Large \sqrt{24} = 2\sqrt{6}\]

OpenStudy (sugarapples):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome